OpenStudy (anonymous):

Hugh made a recipe that yields 5 and 1 over 4 cups. If each serving is 1 over 2 cup, which expression will help Hugh determine the number of servings his recipe will yield? 5 1/4+1/2 would that be correct?

ganeshie8 (ganeshie8):

you need to divide total amount by the each serving amount

ganeshie8 (ganeshie8):

\[\large \dfrac{~~~5\dfrac{1}{4}~~~}{\dfrac{1}{2}}\]

OpenStudy (anonymous):

got it @ganeshie8 thanks

ganeshie8 (ganeshie8):

yw!

OpenStudy (anonymous):

The temperature is changing every day by negative 3 over 16 degree Fahrenheit. What will be the change in the temperature after 4 days? I got 3/4